# ALPCAHUS: Subspace Clustering for Heteroscedastic Data

### v1

@@ V-A2 Clustering Error
ALPCAHUS ($B=1$) with TIPS initialization achieved lower clustering error than KSS with TIPS initialization. For the ensemble methods ($B>1$), both EKSS and ALPCAHUS significantly improved. However, EKSS still had higher clustering error in more heteroscedastic regions. Meanwhile, ALPCAHUS remained very close to the noisy oracle, indicating it more accurately estimated the underlying true labeling. Compared to other methods like TSC and ADSSC, ALPCAHUS generally outperformed them with one exception for ADSSC ($N_{2}=1,\nu_{2}/\nu_{1}=300$). Here, ADSSC performed better than ALPCAHUS when the number of noisy points is roughly equal to the number of good points. Overall, the ensemble version of ALPCAHUS was generally more robust against heteroscedasticity compared to other subspace clustering algorithms. To summarize, we explored the effects of data quality and data quantity on the heteroscedastic subspace clustering estimates in different situations. To our knowledge, this is the first systematic analysis of heteroscedasticity effects on subspace clustering quality in the literature.
@@ V-A3 Effects of Good Data
Additionally, we explored the effects of good data quantity on the subspace clustering problem to understand at what point it becomes advantageous to use ALPCAHUS over KSS methods. We fixed $\nu_{1}=0.1,N_{2}=500$ and vary $N_{1},\nu_{2}$. We report the percentage difference of EKSS clustering error (%) - ALPCAHUS clustering error (%), i.e., $\text{error}_{\text{EKSS}}-\text{error}_{\text{ALPCAHUS}}$, meaning higher values indicate our method is better. Figure 3 shows that ALPCAHUS performed better than EKSS even up to $N_{1}=160$ which represents about 33% of the total data. ALPCAHUS never performed worse than EKSS but the advantage gap narrowed as $N_{1}$ increased. Thus, there is a wide range of conditions for which ALPCAHUS is preferable to homoscedastic methods.
@@ V-A4 Clustering Initialization
Section IV proposed using the TIPS initialization scheme in this heteroscedastic context since the dot product metric used in constructing the affinity matrix in (29) is provably robust to heteroscedastic noise. We fixed points groups $N_{1},N_{2}$ and varied $\nu_{2}$ while keeping $\nu_{1}=0.1$. We included the noisy oracle to establish a realistic performance baseline. Figure 4 shows that TIPS initialization ALPCAHUS $(B=1)$ outperformed random initialization for the non-ensemble ALPCAHUS method $(B=1)$ across the entire heteroscedastic landscape. Thus, TIPS should always be used for the non-ensemble methods for clustering quality. We did not include the ensemble version of ALPCAHUS $(B>1)$ because the ensemble process inherently takes advantage of random initialization to achieve a different labeling each trial.
@@ V-A5 Rank Estimation of Clusters
In Sec. IV, we proposed to use random sign flipping to adaptively find and shrink the subspace dimension of clusters when subspace dimension is unknown. For Fig. 5 we synthetically generated subspaces with true rank $d=6$ and explored how the initial rank parameter affects the ability to estimate the true rank over 100 trials. We compared this approach against the eigengap heuristic by using ALPCAHUS with both adaptive rank schemes and report the estimated rank values from the final clustering. The eigengap approach consistently underestimated the rank regardless of the initial value, except when given the true rank value $d=6$. Our sign flipping approach provided much better rank estimates with much smaller variances between trials.
@@ V-B1 Quasar Flux Data
We investigated quasar spectra data from the Sloan Digital Sky Survey (SDSS) Data Release 16 [39] using its DR16Q quasar catalog [40]. Each quasar has a vector of flux measurements across wavelengths that describes the intensity of observing that particular wavelength. In this dataset, the noise is heteroscedastic across the sample space (quasars) and feature space (wavelength), but we focused on a subset of data that is homoscedastic across wavelengths and heteroscedastic across quasars. The noise for each quasar is known given the measurement devices used for data collection [39]. In Fig. 7, a subset of the spectra data is shown for illustrative purposes to compare the visual differences in data quality. We preprocessed the data (filtering, interpolation, centering, and normalization) based on the supplementary material of [41]. We form clusters in these experiments by considering quasars with different properties, namely, two quasar groups ($K=2$) with different redshift values ($z_{1}=1.0-1.1,z_{2}=2.0-2.1$). Additionally, for the second group, we queried for broad absorption lines (BAL) type quasars. Since we downloaded the data using separate queries, we know which points belong to which cluster group, meaning we can compute clustering error for comparison purposes. We isolated a training set (5000 samples total) to learn any model parameters and used the rest (5000 samples total) for a test dataset. We formed trials by randomly selecting 400 quasar spectra samples per group and performed 50 total trials to report clustering error. For rank estimation, we used the noisy oracle approach and found that there is no improvement to clustering quality for rank values greater than $\hat{d}=3$, so we used this value for any applicable algorithms.
@@
Figure 7: Sample data matrix of quasar flux measurements across wavelengths for each (column-wise) sample.
@@
Figure 8: Experimental results of Indian Pines data for subspace clustering.
@@ V-B2 Indian Pines Data
Additionally, we investigated hyperspectral image (HSI) segmentation data called Indian Pines [42]. This image of size $145\times 145$ contains $D=200$ reflectance bands for each pixel that is around the $0.4-2.5\mu m$ wavelength. HSI data is known to be very noisy due to thermal effects, atmospheric effects, and camera electronics, leading to works that study per-pixel noise estimation in hyperspectral imaging [43]. In total, there are $K=16$ classes composed of alfalfa, corn, grass, wheat, soybean, and others. There is one additional class (background) that is withheld from clustering as commonly done [33]. Therefore, $N=10,249$ instead of $N=145\times 145$. In other works, researchers have found that $\hat{d}=5$ is an appropriate rank parameter since it covers $95\%$ of the cumulative variance [33]. A subset of the data matrix $Y\in\mathbb{R}^{200\times 10249}$ is split into $2,500$ samples to learn model parameters. From this, the learned subspaces are applied on $Y$ to cluster all data so that the heatmaps can be compared against the ground truth. The ground truth image is found in Fig. 8(a). For perspective, randomly guessing would lead to 94% clustering error due to $K=16$ clusters. For consistency, the Hungarian algorithm is again applied to the clustering results in Fig. 8 to make it easier to compare against the ground truth label image in Fig. 8(a).
@@ Methods
We compare against $K$-means in Fig. 8(b) to illustrate the difficulty of the problem. ALPCAHUS results are shown in Fig. 8(d) next to EKSS results in Fig. 8(c). Clustering error is reported for these algorithms along with mean IOU to measure the similarity between the images. ALPCAHUS achieved the lowest clustering error and highest mIOU (53% / 31%) relative to the other approaches such as $K$-means (77% / 16%) and EKSS (64% / 27%). We note that our ALPCAHUS results are similar to [33] even though, in that work, they instead treat the noise as homoscedastic and instead model the signal as heteroscedastic. Further, the EKSS results on this dataset are similar to the homoscedastic PCA approach used in [33]. This indicates some utility in modeling heteroscedasticity in hyperspectral images. Yet, the reflectance bands themselves also appear to be heteroscedastic with some bands being noisier than others [44]. Thus, developing a method that is doubly heteroscedastic with respect to both the samples and features is an interesting direction of future work.
@@ VI Conclusion
This paper proposed ALPCAHUS, a subspace clustering algorithm that can find subspace data clusters whose points contain heteroscedastic noise. For future work, a union of manifolds generalization could possibly be useful. For example, Alzheimer’s disease patients often have different resting state functional MRI activations than cognitively normal individuals [47], so one could consider these different classes to belong to different manifolds in the ambient space. Previous research has shown manifold learning to more correctly model temporal dynamics than subspace modeling in this domain [48]. Another direction of future work could be to consider heteroscedasticity across the feature space. For example, one might be interested in biological sequencing of different species with similar genes where the gene marker counts naturally follow heteroscedastic distributions [49]. These generalizations are nontrivial so it is left for future work. While our implementation of ALPCAHUS benefited from parallelization, it needed more computation time and memory than TSC. Because of this, there is an opportunity to further optimize our code base to remove certain matrix multiplication operations and instead use single vector dot products to reduce memory. Furthermore, since our subspace basis step requires an SVD computation for an initial low-rank estimate, we could use the Krylov-based Lanczos algorithm [50] to further reduce time and memory for each trial leading to more significant improvements in memory usage and time. Overall, ALPCAHUS was more robust towards heteroscedastic noise than other clustering methods, making it easier to identify correct clusterings under this kind of noise model.
@@ VIII-A ALPCAHUS Convergence Experiment
Both upper and lower bound cost function reference values are provided in Fig. 9. The upper bound is generated by first using randomly initialized subspaces as the true subspaces and secondly, using (17) to get cluster labels. From this, the cost function in (15) is compute. Likewise, a lower bound is generated by first using known cluster labels and secondly, estimating the subspaces and noise variances with (16). As observed in Fig. 9, ALPCAHUS converges relatively quickly with only a few iterations. We note that in this figure, the stopping criteria is not implemented. In other words, ALPCAHUS is run for a fixed number of iterations. However, as clearly demonstrated, convergence to a fixed point is achieved earlier for both versions of ALPCAHUS before reaching the maximum number of iterations indicating the usefulness of the stopping criteria to speed up the algorithm.

### v2

@@ V-A2 Clustering Error
ALPCAHUS ($B=1$) with TIPS initialization achieved lower clustering error than KSS with TIPS initialization. For the ensemble methods ($B>1$), both EKSS and ALPCAHUS significantly improved. However, EKSS still had higher clustering error in more heteroscedastic regions. Meanwhile, ALPCAHUS remained very close to the noisy oracle, indicating it more accurately estimated the underlying true labeling. Compared to other methods like TSC and ADSSC, ALPCAHUS generally outperformed them with one exception for ADSSC ($N_{2}=1,\nu_{2}/\nu_{1}=300$). Here, ADSSC performed better than ALPCAHUS when the number of noisy points is roughly equal to the number of good points. Overall, the ensemble version of ALPCAHUS was generally more robust against heteroscedasticity compared to other subspace clustering algorithms. To summarize, the effects of data quality and data quantity on the heteroscedastic subspace clustering estimates is explored in different situations. To our knowledge, this is the first systematic analysis of heteroscedasticity effects on subspace clustering quality in the literature.
@@ V-A3 Effects of Good Data
Additionally, the effects of good data quantity on the subspace clustering problem is explored to understand at what point it becomes advantageous to use ALPCAHUS over KSS methods. The following parameters are fixed $\nu_{1}=0.1,N_{2}=500$ and the following vary $N_{1},\nu_{2}$. The reporting metric is the percentage difference of EKSS clustering error (%) - ALPCAHUS clustering error (%), i.e., $\text{error}_{\text{EKSS}}-\text{error}_{\text{ALPCAHUS}}$, meaning higher values indicate our method is better. Figure 3 shows that ALPCAHUS performed better than EKSS even up to $N_{1}=160$ which represents about 33% of the total data. ALPCAHUS never performed worse than EKSS but the advantage gap narrowed as $N_{1}$ increased. Thus, there is a wide range of conditions for which ALPCAHUS is preferable to homoscedastic methods.
@@ V-A4 Clustering Initialization
Section IV proposed using the TIPS initialization scheme in this heteroscedastic context since the dot product metric used in constructing the affinity matrix in (29) is provably robust to heteroscedastic noise. The following parameters were fixed $N_{1},N_{2}$ and the following varied $\nu_{2}$ while keeping $\nu_{1}=0.1$. The noisy oracle result is included to establish a realistic performance baseline. Figure 4 shows that TIPS initialization ALPCAHUS $(B=1)$ outperformed random initialization for the non-ensemble ALPCAHUS method $(B=1)$ across the entire heteroscedastic landscape. Thus, TIPS should always be used for the non-ensemble methods for clustering quality. The ensemble version of ALPCAHUS $(B>1)$ is not included in these results because the ensemble process inherently takes advantage of random initialization to achieve a different labeling each trial.
@@ V-A5 Rank Estimation of Clusters
In Sec. IV, it is proposed to use random sign flipping to adaptively find and shrink the subspace dimension of clusters when subspace dimension is unknown. For Fig. 5 synthetic subspaces are generated with true rank $d=6$ to explore how the initial rank parameter affects the ability to estimate the true rank over 100 trials. This approach is compared against the eigengap heuristic by using ALPCAHUS with both adaptive rank schemes and report the estimated rank values from the final clustering. The eigengap approach consistently underestimated the rank regardless of the initial value, except when given the true rank value $d=6$. Our sign flipping approach provided much better rank estimates with much smaller variances between trials.
@@ V-B1 Quasar Flux Data
Quasar spectra data from the Sloan Digital Sky Survey (SDSS) Data Release 16 [39] is investigated using its DR16Q quasar catalog [40]. Each quasar has a vector of flux measurements across wavelengths that describes the intensity of observing that particular wavelength. In this dataset, the noise is heteroscedastic across the sample space (quasars) and feature space (wavelength), but for our experiments focused on a subset of data that is homoscedastic across wavelengths and heteroscedastic across quasars. The noise for each quasar is known given the measurement devices used for data collection [39]. In Fig. 7, a subset of the spectra data is shown for illustrative purposes to compare the visual differences in data quality. The preprocessing pipeline for the data included (filtering, interpolation, centering, and normalization) based on the supplementary material of [41]. Clusters are formed in these experiments by considering quasars with different properties, namely, two quasar groups ($K=2$) with different redshift values ($z_{1}=1.0-1.1,z_{2}=2.0-2.1$). Additionally, for the second group, a query for broad absorption lines (BAL) type quasars is done. Since the downloaded data is queried separately, one knows which points belong to which cluster group, meaning one can compute clustering error for comparison purposes. A training set (5000 samples total) is set aside to learn any model parameters and the rest (5000 samples total) is used for a test dataset. Trials are formed by randomly selecting 400 quasar spectra samples per group and running clustering algorithms for 50 total trials. For rank estimation, the noisy oracle approach is used to find that there is no improvement to clustering quality for rank values greater than $\hat{d}=3$, so this value is used for any applicable algorithms that require it.
@@
Figure 7: Sample data matrix of quasar flux measurements across wavelengths for each column-wise sample.
@@
Figure 8: Experimental results of Indian Pines HSI data in a subspace clustering context. Results reported are clustering error and mean IOU (intersection over union) for each algorithm.
@@ V-B2 Indian Pines Data
Hyperspectral image (HSI) segmentation data, called Indian Pines [42], is explored in this section. This image of size $145\times 145$ contains $D=200$ reflectance bands for each pixel that is around the $0.4-2.5\mu m$ wavelength. HSI data is known to be very noisy due to thermal effects, atmospheric effects, and camera electronics, leading to works that study per-pixel noise estimation in hyperspectral imaging [43]. In total, there are $K=16$ classes composed of alfalfa, corn, grass, wheat, soybean, and others. There is one additional class (background) that is withheld from clustering as commonly done [33]. Therefore, $N=10,249$ instead of $N=145\times 145$. In other works, researchers have found that $\hat{d}=5$ is an appropriate rank parameter since it covers $95\%$ of the cumulative variance [33]. A subset of the data matrix $Y\in\mathbb{R}^{200\times 10249}$ is split into $2,500$ samples to learn model parameters. From this, the learned subspaces are applied on $Y$ to cluster all data so that the heatmaps can be compared against the ground truth. The ground truth image is found in Fig. 8(a). For perspective, randomly guessing would lead to 94% clustering error due to $K=16$ clusters. For consistency, the Hungarian algorithm is again applied to the clustering results in Fig. 8 to make it easier to compare against the ground truth label image in Fig. 8(a).
@@ Methods
$K$-means is compared against in Fig. 8(b) to illustrate the difficulty of the problem. ALPCAHUS results are shown in Fig. 8(d) next to EKSS results in Fig. 8(c). Clustering error is reported for these algorithms along with mean IOU to measure the similarity between the images. ALPCAHUS achieved the lowest clustering error and highest mIOU (53% / 31%) relative to the other approaches such as $K$-means (77% / 16%) and EKSS (64% / 27%). We note that our ALPCAHUS results are similar to [33] even though, in that work, they instead treat the noise as homoscedastic and instead model the signal as heteroscedastic. Further, the EKSS results on this dataset are similar to the homoscedastic PCA approach used in [33]. This indicates some utility in modeling heteroscedasticity in hyperspectral images. Yet, the reflectance bands themselves also appear to be heteroscedastic with some bands being noisier than others [44]. Thus, developing a method that is doubly heteroscedastic with respect to both the samples and features is an interesting direction of future work.
@@ VI Conclusion
This paper proposed ALPCAHUS, a subspace clustering algorithm that can find subspace clusters whose samples contain heteroscedastic noise. For future work, a union of manifolds generalization could possibly be useful. For example, Alzheimer’s disease patients often have resting state functional MRI activations different than cognitively normal individuals [47], so one could consider these different classes to belong to different manifolds in the ambient space. Previous research has shown manifold learning to more correctly model temporal dynamics than subspace modeling in this domain [48]. Another direction of future work could be to consider heteroscedasticity across the feature space instead of the sample space. For example, one might be interested in biological sequencing of different species with similar genes where the gene marker counts naturally follow heteroscedastic distributions [49]. These generalizations are nontrivial so it is left for future work.
While our implementation of ALPCAHUS benefited from parallelization, it needed more computation time and memory than TSC. Because of this, there is an opportunity to further optimize our code base to remove certain matrix multiplication operations and instead use single vector dot products to reduce memory. Further, since our subspace basis step requires an SVD computation for an initial low-rank estimate, perhaps the Krylov-based Lanczos algorithm [50] can be used to further reduce time and memory for each trial leading to more significant improvements in memory usage and time. Overall, ALPCAHUS was more robust towards heteroscedastic noise than other clustering methods, making it easier to identify correct clusterings under this kind of heteroscedastic noise model.
@@ VIII-A ALPCAHUS Convergence Experiment
Both upper and lower bound cost function reference values are provided in Fig. 9. The upper bound is generated by first using randomly initialized subspaces as the true subspaces and secondly, using (17) to get cluster labels. From this, the cost function in (15) is computed. Likewise, a lower bound is generated by first using the known cluster labels and secondly, estimating the subspaces and noise variances with (16). As observed in Fig. 9, ALPCAHUS converges relatively quickly with only a few iterations. Note that in this figure, the stopping criteria is not implemented. In other words, ALPCAHUS is run for a fixed number of iterations. However, as clearly demonstrated, convergence to a fixed point is achieved earlier for both versions of ALPCAHUS before reaching the maximum number of iterations. This indicates the usefulness of the stopping criteria to speed up the algorithm in various applications.